OpenStudy (anonymous):

I WILL FAN AND MEDAL!!!!!!!!!! Compare and Contrast: Below are two expressions. Simplify each and then choose the statement that is true. Expression #1(3d)3(d) Expression #2 (3d2)2 The exponents in Expression #1 are greater than the exponents of Expression #2. The exponents on Expression #2 are greater than the exponents of Expression #1. The exponents of Expression #1 are the same as the exponents of Expression #2. The relationship cannot be determined with the given information.

OpenStudy (mathstudent55):

Expression 1: \((3d^3)(d) \) Expression 2: \( (3d^2)^2 \) Did I understand both expressions correctly?

OpenStudy (anonymous):

Yup you did @mathstudent55

OpenStudy (mathstudent55):

Let's simplify each one: 1. \( (3d^3)(d) = 3d^3 \times d^1 = 3d^4\) 2. \((3d^2)^2 = 3^2 \times (d^2)^2 = 9 d^{2 \times 2} = 9 d^4 \)

OpenStudy (anonymous):

Ok yes correct i see where your going

OpenStudy (mathstudent55):

Wait. Isn't Expression 1: \( (3d)^3 d \) ?

OpenStudy (anonymous):

Yeah expression 1 is that

OpenStudy (mathstudent55):

I originally put the parentheses in the wrong place. 1. \( (3d)^3 d = 3^3 d^3 d = 27 d^4 \)

OpenStudy (anonymous):

Lol Ok I see what you mean

OpenStudy (anonymous):

So what does that mean @mathstudent55

OpenStudy (mathstudent55):

I'm not sure what they mean with the choices. In the end, the exponent of d is 4 in either expression. If that's what they mean then the exponents are equal. That means choice C. On the other hand, Expression 2 ends up with 9 which is \(3^2 \) and Expression 1 ends up with 27 which is \( 3^3 \), so here exponent 3 is greater than exponent 2. Since the numbers can be written as 27 and 9 with no exponents, maybe they mean choice C, but I am not sure because I don't quite understand what they are asking.